### v1

# https://xiaojiu-z.github.io/Stable-Makeup.github.io/

Abstract
Current makeup transfer methods are limited to simple makeup styles, making them difficult to apply in real-world scenarios. In this paper, we introduce Stable-Makeup, a novel diffusion-based makeup transfer method capable of robustly transferring a wide range of real-world makeup, onto user-provided faces. Stable-Makeup is based on a pre-trained diffusion model and utilizes a Detail-Preserving (D-P) makeup encoder to encode makeup details. It also employs content and structural control modules to preserve the content and structural information of the source image. With the aid of our newly added makeup cross-attention layers in U-Net, we can accurately transfer the detailed makeup to the corresponding position in the source image. After content-structure decoupling training, Stable-Makeup can maintain content and the facial structure of the source image. Moreover, our method has demonstrated strong robustness and generalizability, making it applicable to various tasks such as cross-domain makeup transfer, makeup-guided text-to-image generation and so on. Extensive experiments have demonstrated that our approach delivers state-of-the-art (SOTA) results among existing makeup transfer methods and exhibits a highly promising with broad potential applications in various related fields.
Keywords:
Makeup Transfer, D-P Makeup Encoder, Diffusion
Figure 1: Our proposed framework, Stable-Makeup, is a novel diffusion-based method for makeup transfer that can robustly transfer a diverse range of real-world makeup styles, from light to extremely heavy makeup.
*** Work done during internship at Xiaohongshu Inc.
$\dagger$ Corresponding author
1 Introduction
As a significant computer vision task, makeup transfer has a wide range of applications, such as in the beauty industry or in virtual try-on systems, and enables the enhancement, modification, and transformation of facial features, achieving the desired effects of beautification, embellishment, and deformation. However, despite its conceptual straightforwardness, makeup transfer poses significant challenges when aiming for a seamless and authentic transformation across a wide range of makeup intensities and styles.
From the perspective of the existing technical route, current makeup transfer techniques [21, 17, 7, 26, 47, 12, 22, 43, 50, 40, 19, 51] rely heavily on Generative Adversarial Network (GAN) based approaches which have demonstrated potential in digital cosmetic applications. However, these approaches fall short when confronted with the diversity of real-world makeup styles, especially when translating high-detailed and creative cosmetics, such as those found in cosplay or movie character imitations, onto real faces. This limitation not only restricts their applicability but also hinders their effectiveness in accurately capturing the essence of personalized and intricate makeup designs. Recognizing this gap, our research introduces Stable-Makeup, a novel approach leveraging diffusion-based methodologies to transcend the boundaries of existing makeup transfer methods.
From a data perspective, existing makeup datasets lack diversity and cannot accommodate real-world makeup transfer. To address this data shortage, we propose an automatic data construction pipeline that employs large language model and generative model to edit real human face images and create paired before-and-after makeup images. Ultimately, our dataset comprises 20k image pairs, covering a wide range of makeup styles from light to heavy makeup. This dataset has the potential to facilitate research in the makeup transfer field and enable the development of more robust and accurate models.
Our Stable-Makeup, which is built on a pre-trained diffusion model, and comprises three key components: the Detail-Preserving Makeup Encoder, Makeup Cross-attention Layers and the Content and Structural Control Modules. To preserve makeup details, we employ a multi-layer strategy in our D-P makeup encoder to encode the makeup reference image into multi-scale and spatial-aware detail makeup embeddings. The content control module is designed to maintain pixel-level content consistency with source image. The structural control module is utilized to introduce facial structure, improving the consistency between the generated image and the facial structure of the source image. To achieve semantic alignment between the intermediate features of the Unet-encoded source image and the detail makeup embeddings, we extended the U-Net architecture by incorporating a makeup branch composed of cross-attention layers. Through our proposed content and structural decoupling training strategy, we can further maintain the facial structure of the source image. As exhibited in Fig. 1, By integrating a meticulously curated dataset alongside a comprehensive pipeline from data acquisition to model training, we lay the foundation for a robust framework capable of handling the nuanced dynamics of makeup transfer with unprecedented precision and adaptability.
In summary, our contributions are:
  • In this paper, we present Stable-Makeup, the first diffusion-based makeup transfer method to the best of our knowledge. Our experimental results demonstrate state-of-the-art performance over existing makeup transfer methods. Moreover, Stable-Makeup demonstrates high robustness and generalizability, making it applicable to various tasks.
  • To ensure the accuracy and consistency of makeup transfer, we propose a Detail-Preserving makeup encoder in conjunction with makeup cross-attention layers that align the semantic correspondence between detailed makeup features and human face intermediate features in U-Net.
  • We designed an automatic pipeline to create a high-quality, diverse paired dataset of makeup before and after images. After content-structure decoupling training on this dataset, we can further maintain the content and structure of the source image.
2 Related Works
2.1 Facial Makeup Transfer
Traditional makeup transfer methods employ image processing techniques such as facial landmark extraction and detection[49, 41]. Meanwhile, contemporary deep learning-based makeup transfer methods demonstrate high robustness and generalizability to diverse makeup styles, Generative Adversarial Networks (GANs)[58, 5, 21, 17, 7, 26, 47, 12, 22, 43, 50, 40, 19, 15, 51] as an example, have been extensively applied in facial makeup synthesis tasks.
In the domain of image transfer, Beauty-GAN[21] driven by pixel-level Histogram Matching and utilizes an array of loss functions to train the primary network. PSGAN[17] addresses the challenge of transferring makeup between images with significant facial expression variations and focuses on specific facial regions for transfer. To expand the scope of makeup transfer tasks beyond mere color transfer in specific facial areas, CPM[26] introduces pattern addition into the makeup transfer process. Furthermore, SCGAN[7] leverages a Part-specific Style Encoder to decouple makeup styles on a component-wise basis. Additionally, RamGAN[47] focuses on component consistency by integrating a region-attentive morphing module.
The inadequacies of GAN-based models in addressing makeup transfer problems under real-world makeup settings have necessitated the development of new techniques. The primary objective of this paper is to overcome these limitations by utilizing diffusion-based methods to achieve robust and high-quality output for extreme style makeup.
2.2 Diffusion Models
Recently, diffusion models have ushered in significant advancements in multi-modal image generation tasks, encompassing text-to-image generation[30, 27, 34, 31, 35], image editing[6, 20, 9, 3, 48, 24, 2, 56, 42], controllable generation[53, 25, 57, 23] and subject-driven generation[55, 14, 1, 16, 10, 33, 32]. Methods based on diffusion paradigm, along with their subsequent adaptations, have demonstrated state-of-the-art results in various image generation tasks.
The diffusion model was initially introduced based on the physical inverse diffuse process [37]. Subsequently, DDPM[13] demonstrated that this approach is viable for image generation by denoising a Gaussian-noised image. DDIM[38] introduced a faster non-Markovian sampling process and an implicit modeling of the denoising step, both of which significantly enhance the model’s efficiency. Latent Diffusion Models[31] utilize an Autoencoder to encode images into latent embeddings, effectively increasing computational efficiency while preserving the essential features of the input image. In recent years, there have been significant advancements in diffusion-based text-to-image generation. For instance, Imagen[34] and Stable Diffusion[31] have trained diffusion models on larger datasets, making them the mainstream for generating large images. DeepFloyd IF[35] has utilized a triple-cascade diffusion model, which has significantly enhanced the text-to-image generation capability, even generating correct fonts. Stable Diffusion XL[27], a two-stage cascade diffusion model, which has greatly improved the generation of high-frequency details, small object features, and overall image color. As for controllable generation, ControlNet[54] employs a bypass structure featuring a trainable copy network as the input mechanism for control conditions, effectively integrating a variety of control conditions such as Canny Edge, Depth, and Human Pose, etc. As for subject generation, DreamBooth[32] Dreambooth trains a separate Unet for each subject, resulting in excellent consistency of character. However, the high computational cost limits its practical application. Meanwhile, diffusion based image editing methods have also demonstrated remarkable progress. For example, Ledits[42] introduces a training-free image manipulation method, utilizing an inversion approach for processing.
Due to the remarkable advancements of diffusion methods in generating high-quality images, benefiting from supervised training on large-scale datasets, they have become the mainstream method in image generation. Therefore, in this paper, we have leveraged the advantages provided by the diffusion model to the fullest extent in the task of makeup transfer. Our approach overcomes the limitations of GAN-based models and achieves robust and high-quality output, particularly for extreme style makeup.
3 Methodology
In this part, we introduce the detailed design of our Stable-Makeup in Section 3.1. Then, the specifics of our training and inference process are detailed in Section 3.2. Preliminaries about diffusion models [31], Controlnet [54] and CLIP [29] can be found in supplementary.
3.1 Stable-Makeup
3.1.1 Overview.
The overall methodology is illustrated in Fig. 2. Formally, the target of Stable-Makeup is to transfer the reference makeup from a given image $\mathit{I_{m}}$ onto a user-provided source image $\mathit{I_{s}}$, resulting in a target image $\mathit{I_{t}}$ that exhibits the desired makeup style. Therefore, Stable-Makeup necessitates the simultaneous transfer of fine-grained makeup information and preservation of structural information. To achieve this, Stable-Makeup first utilizes a Detail-Preserving Makeup Encoder to extract the multi-scale and spatial-aware features of the reference makeup as detailed makeup embeddings $\mathit{E_{m}}$. Next, content encoder and structural encoder are employed to encode the source image as content features $\mathit{F_{c}}$ and the facial structure control image $\mathit{I_{c}}$ (obtained from an off-the-shelf face parsing method [28].) as structure features $\mathit{F_{s}}$, respectively. These two features are then incorporated together and fed into the U-Net. The detailed makeup embeddings $\mathit{E_{m}}$ along with the content features $\mathit{F_{c}}$ and structure features $\mathit{F_{s}}$ are integrated into the diffusion U-Net as conditions to guide the image generation process. During target image $\mathit{I_{t}}$ generation, Stable-Makeup uses newly added cross-attention layers to implicitly align the multi-scale detailed makeup embeddings $\mathit{E_{m}}$ with the intermediate feature maps of facial region in source image at different layers in U-Net. The alignment ensures that the generated target image contains both the detailed information of the reference makeup and maintains the content and structure of the source image $\mathit{I_{s}}$. With our Content and Structure Decoupling Training, Stable-Makeup can further maintain the content and facial structure of the source image.
Figure 2: Overall schematics of our method. Given a source image $\mathit{I_{s}}$, a reference makeup image $\mathit{I_{m}}$ and an obtained facial structure control image $\mathit{I_{c}}$, Stable-Makeup utilizes D-P makeup encoder to encode $\mathit{I_{m}}$. Content and structural encoders are used to encode $\mathit{I_{s}}$ and $\mathit{I_{c}}$ respectively. With the aid of the makeup cross-attention layers, Stable-Makeup aligns the facial regions of $\mathit{I_{s}}$ and $\mathit{I_{m}}$, enabling successful transfers the intricate makeup details. After content-structure decoupling training, Stable-Makeup further maintains content and structure of $\mathit{I_{s}}$.
3.1.2 Detail-Preserving Makeup Encoder.
Makeup images often involve intricate details such as eyelashes, eyebrows, and fine lines. These details are often subtle and precise, making it challenging to maintain their quality during makeup transfer. To address this issue, we have carefully designed a Detail-Preserving Makeup Encoder to capture the details of the makeup.
[…]
Notably, some previous works[52, 46, 11] utilize linear layer mapping of CLIP features. This design cannot preserve the spatial information between input features. Different from them, we utilize a self-attention layer to map the multi-layer CLIP embeddings (256-dimensional patch features and 1-dimensional CLS features) of the reference makeup image and concatenate on feature dimension. This approach effectively preserves the spatial, multi-scale, and global information of the makeup image, thereby facilitating improved alignment of intermediate related features within the U-Net architecture. Consequently, it yields more precise preservation of fine-grained details and enhanced transfer outcomes. Combining both spatial and scale feature maps, our D-P makeup encoder yields the most complete information for makeup. Furthermore, we provide additional comparative analysis in supplementary to demonstrate the superiority of our method.
3.1.3 Makeup Cross-attention Layers.
The goal of the makeup cross-attention layers is to integrate multi-scale detailed makeup embeddings into the U-Net and align them appropriately with the intermediate feature maps of facial region in the source image, enabling the model to accurately transfer the makeup onto the facial region in the source image. Specifically, as shown in the figure 2, in each transformer block of the U-Net, we retain the original text cross-attention and add Makeup Cross-attention Layers. The multi-scale detailed makeup embeddings $\mathit{E_{m}}$ are input to the makeup cross-attention layers, serves as the K(key) and V(value) features. The makeup cross-attention layers and the text cross-attention layers in the original U-Net share the Q feature. With the help of the cross-attention mechanism, the intermediate facial features of source image and makeup image are well-aligned in the stable diffusion U-Net. Finally, we simply add the output of makeup cross-attention to the output of text cross-attention.
Our design of adding the output of the cross-attention between the makeup and the original text preserves the original text-guided generation capability. In the context of the classic makeup transfer task, the influence of text is not considered, only the makeup cross-attention is employed during both training and inference. However, our design opens up a new direction, namely makeup-guided text-to-image generation, as illustrated in Figure 6. In this task, the reference makeup image is employed as a guiding condition in conjunction with the original text prompt to influence the pre-trained U-Net architecture, and the makeup consistency is maintained during the text-to-image generation process.
3.1.4 Content and Structural Control Modules.
To maintain the structural consistency of the source image during makeup transfer, we utilize Content and Structural Control Modules, which consists of two adapted ControlNets serving as encoders. The target of the content encoder is to maintain content consistency and consistency in non-facial areas of the original image, and it uses the ControlNet to encode the source image. To control facial structures during the makeup transfer process, we added an extra structural encoder. In order to increase spatial constraints and introduce more facial information, such as facial shape or mouth closure, we utilized dense lines of varying colors that were drawn based on the different locations of facial key points. These lines served as conditions for the makeup transfer process, enabling us to achieve great control over the resulting facial structures. We use the text embedding of empty text as the cross-attention input for the two ControlNet encoders. Finally, we directly add the outputs of the content encoder and structural encoder to the Stable Diffusion model.
3.2 Model Training and Inference
3.2.1 Makeup-real Dataset.
As depicted in Fig. 3, Ours pipeline to create a large-scale makeup pairing dataset (Makeup-real dataset) by leveraging the capabilities of LLM and pre-trained diffusion involves three main steps: Firstly, we utilize GPT4 to generate 1000 prompts using the template "Make it {} makeup" to cover a wide range of makeup styles. After filtering out duplicate words, we obtain 300 different prompts. Secondly, we sample 20,000 faces from the FFHQ[18] dataset as non-makeup images and apply makeup to these images using existing text-based real image editing algorithm [42]. For each face image, we randomly select a makeup style from the 300 prompts obtained in the first step, resulting in 20,000 face images with consistent identities and makeup styles. Finally, to prevent the unwanted effects of editing non-face regions during the edit process, we use a face segmentation model to separate the face region from the background and other non-face regions. We composite the non-face regions of the original face images onto the edited face images to obtain the final makeup pairing dataset. [Uncaptioned image] Figure 3: The procedure for training data generation. Finally, our dataset includes 20,000 image pairs covering a wide range of makeup styles from light to heavy makeup, represents a significant improvement over previous datasets in terms of diversity and quantity. In our supplementary materials, we also present additional visual results for the alternative methods in Step 2.
3.2.2 Content-Structure Decoupling Training.
The current diffusion-based real image editing faces a conflict between maintaining the original layout and editability. This can result in subtle changes in facial structures that are visually perceptible before and after editing. As shown in Fig. 5 (b) fourth column, if left unaddressed, this can lead to inconsistent facial structures between the generated images and the source images after makeup transfer, which is not aligned with the task objective. To tackle this challenge, we propose a training strategy that decouples content and structure to eliminate the impact of inconsistent facial Structures between training data pairs and ensure the generated images have consistent facial structures with the source images.
Specifically, during training, $\mathit{I_{s}}$ and $\mathit{I_{c}}$ obtained from the target image are fed into the content encoder and structural encoder, respectively, while the augmented $\mathit{I_{t}}$ is used as the input for the makeup encoder. This implies that the facial structures of the final generated target image are determined by the structural encoder, while the content is determined by the content encoder. The pre-trained U-Net is used to predict and remove the noise that needs to be eliminated for reconstructing the target image. By decoupling content and structure during training, we can control and adjust the content and structure separately, improving the accuracy and consistency of the generated images. During inference, $\mathit{I_{s}}$ and $\mathit{I_{c}}$ obtained from the source image are input to content encoder and structural encoder respectively.
3.2.3 Augmentation.
In our training pipeline, we have used a variety of augmentations, which are crucial for adapting to real-world scenarios and achieving successful makeup transfer. These augmentations can be divided into two categories: makeup augmentations and source augmentations. Makeup augmentations involves structural augmentations such as face warping and affine transformations applied to the input of the makeup encoder. The purpose of these augmentations is to eliminate Content information from the reference makeup image and to disrupt the pixel alignment between the reference makeup image and the source image. Source augmentations are designed to make our model adaptable to varying facial sizes and poses in the source images in real-world scenarios. These augmentations include synchronized affine transformations applied to both the source image and the target image.
3.2.4 Loss function.
Our loss function is similar to the original Stable Diffusion training objective function and can be mathematically represented as:
$L(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\mathbf{c_{i}},\mathbf{c_{e}},\mathbf{c_{m}}\right)\right\|_{2}^{2}\right],$ (1)
[…]
4 Experiments
4.1 Experimental Setup
We employed Stable Diffusion V1-5 as the pre-trained diffusion model and trained it on the Makeup-real Dataset, which consists of 20k makeup-non-makeup pairs and each pair with the same identity. We focus on real-world conditions and utilize extremely aggressive augmentation, as described in Section 3.2. While training, the parameters of the D-P makeup encoder, makeup cross-attention layers, content encoder and structural encoder are updated, while the pre-trained text-to-image model remains frozen. The model underwent 100,000 iterations of training on 8 H800 GPUs, with a batch size of 16 per GPU and a learning rate of 5e-5. Inference was performed using DDIM as the sampler, with a step size of 30 and a guidance scale set to 1.5.
4.2 Evaluation Benchmarks
  • CPM-real[26] dataset: It is very diverse in terms of makeup styles, containing both color and pattern makeups. The degree of makeup can vary from light to heavy, from color-oriented to pattern-driven.
  • Makeup-wild[17] dataset: It contains facial images with various poses and expressions as well as complex backgrounds to test methods in the real-world environment.
4.3 Evaluation Metrics
To provide a thorough and objective assessment of the performance of each algorithm in different aspects of makeup transfer, we employed the following metrics: CLIP-I[29] and DINO-I[4] were utilized to assess makeup transfer performance. After makeup transfer, the content and structure information of the input image should not be changed. So we use SSIM [45] and L2-M to evaluate the degree of content and structural maintenance.
Notably, CLIP-I and DINO-I are metrics used to calculate the cosine similarity between the image embeddings of the target image and the makeup image. They utilize the CLIP image encoder or DINO backbone to extract image features. L2-M metric utilizes the facial segmentation method to obtain the background images of the target and the source image, calculate the square of the difference, and then average it based on the number of pixels.
4.4 Comparison Methods
We performed comprehensive comparisons with the most representative makeup transfer algorithms currently available, including BeautyGAN[21], CPM[26], PSGAN[17], EleGANt[51], SCGAN[7], SPMT[59] and SSAT[40], and conduct comparisons on M-Wild dataset and CPM-real dataset respectively.
4.5 Experiment Results
4.5.1 Qualitative Comparison.
Figure 4: Qualitative comparison of different methods. Our results outperform other methods in terms of makeup detail transfer, ranging from light makeup to personalized heavy makeup.
Fig. 4 provides a visual comparison of our proposed Stable-Makeup method and other makeup transfer methods. To further evaluate the performance of these methods on various makeup styles, we used some popular makeup images from social media as reference makeup in addition to the test set. Overall, most methods maintained good consistency in content and structure, with only CPM introducing some artifacts during makeup transfer. In terms of makeup transfer capability, our method outperformed other methods by a significant margin. Other methods were only effective in transferring simple lip and eye shadow colors, and were only effective for light makeup. In contrast, our method not only performed well for light makeup, but also effectively aligned and transferred heavy makeup to the corresponding positions in the source image, while preserving makeup details and corresponding color distributions on the facial regions of the makeup image.
4.5.2 Quantitative Comparison.
Table 1: Quantitative comparison of different methods. Metrics that are bold and underlined represent methods that rank 1st and 2nd, respectively.
Method CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$
(M-wild dataset) (CPM-real dataset)
BeautyGAN 0.818 0.927 0.878 48.775 0.775 0.913 0.858 53.323
CPM 0.815 0.923 0.862 108.374 0.793 0.921 0.837 114.288
PSGAN 0.739 0.929 0.850 48.816 0.693 0.915 0.839 50.137
EleGANt 0.827 0.928 0.891 46.939 0.787 0.917 0.872 48.995
SCGAN 0.816 0.927 0.878 47.595 0.776 0.915 0.851 50.062
SPMT 0.834 0.930 0.911 31.671 0.778 0.918 0.905 32.323
SSAT 0.823 0.931 0.901 34.512 0.773 0.910 0.881 35.205
Ours 0.860 0.935 0.906 31.483 0.825 0.928 0.883 32.255
To test the ability of methods in the real-world environment, We use the non-makeup 334 images of the Makeup-Wild(M-Wild) dataset as source images, and randomly selected 334 images from the makeup images in the M-Wild dataset and the CPM-real dataset, respectively, as our reference makeup images.
As shown in Table 1, It is evident that Stable-Makeup outperforms all other makeup transfer methods in terms of the DINO-I and CLIP-I metrics, indicating that our method has superior makeup transfer capabilities compared to other methods. Additionally, the SSIM and L2-M metrics achieved good scores when compared with other methods, indicating that our method is able to maintain good content and structure information in the transferred images. It should be noted that SSIM metrics may not accurately reflect the makeup transfer performance, as they may produce good scores when the transferred result is identical to the source image, indicating that the transfer algorithm did not work. In line with previous studies[26, 47, 59], we also listed it as a reference in our evaluation.
4.5.3 User Study.
Table 2: User study. Metrics that are bold represent method that rank 1st. Our method achieves the highest average score.
Method BeautyGAN CPM PSGAN EleGANt SCGAN SPMT SSAT Ours
Average Score 3.01 3.13 2.95 3.62 2.90 3.35 2.85 4.58
We conducted a user study to compare our method with other methods perceptually. We randomly select 10 source images and 10 reference images from both the M-Wild dataset and CPM-real dataset. For each evaluation, each user will see one source image with a reference makeup image, and one image generated by each method. 60 evaluators were asked to provide a comprehensive score for each generated image based on its makeup consistency, content and structure consistency, using a scale of 1 (worst) to 5 (best). The results of average scores in Table 2 indicate that our method outperforms the comparison methods. Our method achieves the highest average score, which clearly indicates that our method outperforms the comparison methods.
4.5.4 Ablation Study.
Figure 5: Qualitative ablation results. Figure (a) explores the classifier-free guidance scale. Figure (b) presents the ablation study of different training and inference settings.
Ablation Setups CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$
(CPM-real dataset)
W.O. structure input 0.776 0.923 0.772 103.029
W.O. source aug 0.816 0.921 0.874 153.463
W.O decoupling training 0.806 0.927 0.791 32.856
W.O multi-layer 0.803 0.911 0.880 34.861
Ours full 0.825 0.928 0.883 32.255
Table 3: Quantitative ablation results on CPM-real dataset.
Firstly, we explored the effect of different parameters for the classifier-free guidance scale (GS) under makeup conditions in figure 5 (a). Next, in figure 5 (b), we conducted ablation experiments by systematically removing each module using the same experimental setup. As shown in figure 5 (a), it was found that a larger GS led to a greater degree of makeup transfer, while our method maintained good content and structural consistency of the source image under different guidance parameters. In figure 5 (b), it demonstrated that without the structural encoder input during inference (first column), the inconsistency in facial structure was observed, such as changes in the shape of eyebrows and mouth in the first column of the generated images. Without makeup augmentation during training (second column), the model does not possess the capability to transfer makeup, but rather, it solely transfers the facial region of the makeup image to the target image. Removing source augmentation during training (third column), the alignment and transferability of the model were weakened, and the background color was changed. When removing structural encoder during training (fourth column, W.O. decoupling training), the facial structure also could not be well preserved. Without the multi-layer strategy for the makeup encoder (fifth column), the transfer of makeup details was incomplete. When our method was fully implemented, it achieved a complete transfer of makeup details while maintaining good consistency of the content and structure of the source image.
Quantitative ablation results are shown in Table 3. We performed quantitative ablation experiments using the same experimental setup on CPM-real dataset as Sec. 4.5. We observed that our method achieved optimal performance when all modules were retained, while removing the structural encoder input during inference (first row) or training (third row) led to a significant decrease in content and structure consistency. Removing the multi-layer strategy during training resulted in a significant decrease in makeup transfer ability, and the absence of source enhancement led to a significant decrease in background consistency. We exclude the setting without makeup augmentations, as shown in Figure 5 (b) second column, where the model can only perform face swapping and lacks the ability to perform makeup transfer. This could significantly affect the calculation of our makeup transfer metrics.
Figure 6: The visual results of Stable-Makeup demonstrate its robustness and generalizability for various applications.
4.6 More Applications
As illustrated in the Fig. 6, our Stable-Makeup not only achieves unprecedented makeup transfer effects on real faces, but also enables various applications that were previously unattainable with traditional makeup transfer methods. For instance, our method can perform makeup transfer on cross-domain faces and guide text-to-image generation with reference makeup conditions to create creative artistic images. Our method can also perform makeup transfer on videos by utilizing effective multi-frame concatenation. Moreover, our approach supports other domain’s reference makeup, such as animal-inspired subjects, animated characters and so on. providing a broader range of creative makeup options for real human faces.
5 Conclusion
In this paper, we propose the Stable-Makeup, a diffusion-based framework aiming for real-world makeup transfer. This method has made a significant breakthrough in the field of makeup transfer by achieving effects that were previously unattainable. The Stable-Makeup consists of detail-preserving makeup encoder, makeup cross-attention layers, content and structural control modules. The detail-preserving makeup encoder captures fine details of makeup, while two encoder structures are employed for the content and structural control modules to encode content and structure of source image. The makeup cross-attention layers effectively aligns facial region between reference makeup and source image. Through the content and structure decoupling training, we are able to further maintain content and structure consistency. In addition, we proposed a automatic pipeline to create a variety of makeup pairing data for training. Our proposed method has surpassed all existing makeup transfer algorithms and has expanded the boundaries of the makeup transfer field. This breakthrough has significant implications for the practical application of makeup transfer technology in various fields, such as cosmetics, entertainment, and fashion.
Supplementary Materials
F Preliminaries
F.0.1 Diffusion models.
Diffusion Model (DM) [13, 39] belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to target data distribution $\mathbf{x_{0}}$ by means of an iterative denoising procedure. Latent Diffusion Model (LDM) [31] is proposed to model image representations in autoencoder’s latent space. LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions. The LDM loss is:
$L_{LDM}(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\bm{\tau_{\theta}}(\mathbf{c})\right)\right\|_{2}^{2}\right],$ (2)
[…]
Stable Diffusion (SD) is a widely adopted text-to-image diffusion model based on LDM. Compared to LDM, SD is trained on a large LAION [36] dataset and replaces BERT with the pre-trained CLIP [29] text encoder.
F.0.2 ControlNet.
ControlNet [54] is one of the most popular control modules in current diffusion models. It receives inputs from various modalities as spatial control conditions, guiding the diffusion model to generate desired images, thereby achieving controllable generation. ControlNet copies an identical U-Net structure as trainable parameters and locks the original U-Net parameters, The trainable copy is connected to the locked model with zero convolution layers, denoted $\mathcal{Z}(\cdot;\cdot)$. The complete ControlNet can represented as:
$\boldsymbol{y}_{\mathrm{c}}=\mathcal{F}(\boldsymbol{x};\Theta)+\mathcal{Z}\left(\mathcal{F}\left(\boldsymbol{x}+\mathcal{Z}\left(\boldsymbol{c};\Theta_{\mathrm{z}1}\right);\Theta_{\mathrm{c}}\right);\Theta_{\mathrm{z}2}\right)$ (3)
Where $\Theta_{\mathrm{z}1}$ and $\Theta_{\mathrm{z}2}$ represent two different zero conv layers’ parameters respectively. $\mathcal{F}$ represent diffusion U-Net, $\boldsymbol{x}$ represent image latent, $\Theta$ is the frozen weight of the U-Net and $\Theta_{\mathrm{c}}$ is the trainable copy weight of the U-Net.
F.0.3 CLIP.
CLIP [29] consists of two integral components: an image encoder, represented as $F(x)$, and a text encoder, represented as $G(t)$. The image encoder, $F(x)$, transforms an image $x$ with dimensions $\mathbb{R}^{3\times H\times W}$ (height $H$ and width $W$) into a $d$-dimensional image feature $f_{x}$ with dimensions $\mathbb{R}^{N\times d}$, where $N$ is the number of divided patches. On the other hand, the text encoder, $G(t)$, creates a $d$-dimensional text representation gt with dimensions $\mathbb{R}^{M\times d}$ from natural language text $t$, where $M$ is the number of text prompts. After training on a contrastive loss function, CLIP can be applied directly for zero-shot image recognition without the need for fine-tuning the entire model.
G Further Exploration of Data Construction
Currently, diffusion-based methods have demonstrated remarkable performance in terms of fidelity and diversity, making them a crucial tool for achieving high-quality and diverse data in step 2 of our data construction process. In addition to the inversion-based image editing methods we employed, we could also choose to use existing identity customization methods, such as InstantID[44], which do not require test-time fine-tuning. However, as shown in figure 7, when using these diffusion-based methods for editing facial images, visible changes in facial structure may occur during makeup generation guided by text. For example, the mouth shape of these generated images cannot be exactly the same. Given the current limitations of diffusion-based methods in terms of fine-grained control, our proposed content and structure decoupling strategy is particularly important for maintaining the facial expression and structure.
Figure 7: Visual results of different data construct pipeline.
H Designing Choice of Detail-Preserving Makeup Encoder
In this section, we conducted a comparative study between different mapping structure setting under the same training conditions to demonstrate that mapping without preserving spatial information struggling to align and preserving makeup details at a fine-grained level, and to validate the effectiveness of our proposed method. We first introduced our experimental settings in Section 8. Subsequently, we conduct visual comparison in Sec. H.2 to demonstrate the pivotal role of our proposed method in preserving and aligning the makeup details. Additionally, in Sec. H.3 we further demonstrated the effectiveness of our proposed method in preserving and aligning the makeup details by visualizing the attention maps of the makeup cross-attention layers in U-Net.
H.1 Experimental Setup
Figure 8: The structural composition of the five distinct mapping schemes. We omitted the drawing of the global vector (cls token from CLIP) for the sake of brevity and clarity.
As shown in Fig. 8, we conducted training on five sets of mapping schemes.
  • Linear mapping: We utilized linear layers to project the spatial dimension to a smaller dimension of (1, 12, 1024), where the original dimension was (1, 257$\times$12, 1024).
  • Spatial linear mapping: We utilized linear layers to project the feature dimension to a smaller dimension without affecting the spatial features dimension, where the original dimension was (1, 257$\times$12, 1024), and the projected dimension was (1, 257$\times$12, 768).
  • Self-Attention mapping: We utilized our proposed self-attention mapping in section 3.1.2. The projected dimension is the same as the original dimension (1, 257$\times$12, 1024).
  • IP-Adapter: We replaced our D-P makeup encoder with an IP-Adapter that projects the pooled feature of CLIP (1, 1, 768) to a higher dimension features (1, 4, 768) using linear layer and layer norm.
  • IP-Adapter Plus: We replaced our D-P makeup encoder with an IP-Adapter Plus, which use a small query network to learn features. Specifically, 16 learnable tokens are defined to extract information from the grid features using a lightweight transformer model. The token features from the query network serve as input to the cross-attention layers. Where the original dimension was (1, 257, 1024), and the final dimension was (1, 16, 768).
All of these models were trained on the Makeup-real dataset using OpenCLIP ViT-L/14 as the CLIP model and StableDiffusion V1-5 as the pre-trained diffusion model. The training parameters were kept consistent, with a learning rate of 5e-5, batch size of 8, and training on 2 H800s. It is noteworthy that only spatial linear mapping and our self-attention mapping preserve the spatial information of the makeup image features among the five mapping settings, while the others introduce varying degrees of spatial distortion to the makeup image features.
Figure 9: Visual comparison results under the same training step but with different mapping settings.
H.2 Comparison
Under the same training conditions, the makeup transfer results at different training steps are presented in figure 9. Overall, our self-attention mapping achieves the most precise makeup transfer and the smallest number of model training convergence steps. Comparing the results under different settings, it is observed that spatial linear mapping, IP-Adapter, and IP-Adapter plus all maintain the content and structure consistency of the source image, but they fail to align and transfer the details of the reference makeup image to the facial region of source image. Both spatial linear mapping and our self-attention mapping can align and transfer the makeup, but our method outperforms spatial linear mapping in terms of the precision of makeup transfer. For instance, our method completely preserves the red graffiti on the forehead of the clown (in the fifth line), while spatial linear mapping fails to do so (in the third line). It is noteworthy that our method achieves good transfer performance at the 20,000th training step, while spatial linear mapping struggles to preserve the makeup details. These experimental results effectively demonstrate the importance of preserving spatial information for preserving fine makeup details.
H.3 Visualization of Attention maps
Figure 10: Visualization of attention maps with different mapping settings at the 50,000th training step.
In order to conduct a more comprehensive comparison between spatial linear mapping and self-attention mapping, we visualized the attention maps of the makeup cross-attention layers under both configurations at the 50,000th training step in this experiment. Specifically, we extract and aggregate the attention matrices calculated from the patch features of the source facial region (Query) and the multi-scale makeup embeddings (Key-Value), based on the index of the facial position, and visualize them at different layers, as shown in Figure 10 (a). From the attention maps, we can see that the self-attention layer can effectively weight the features, resulting in high response for the facial patch features in between the source image and the makeup image, while low response between the other positions in the makeup image. However, when we replace the self-attention mapping with spatial linear mapping, the facial region of the source image also exhibit high response to the background features, as shown in the attention maps of layer1-4 in Figure 10 (b). These findings highlight the advantages of self-attention mapping in achieving more precise and effective feature weighting, which is crucial for achieving high-quality makeup transfer results.
I More Results
In this section, we present a range of additional results that demonstrate the robustness and superiority of our approach. Specifically, as shown in Fig. 11 and Fig. 12, we showcase the effectiveness of our method in makeup transfer on realistic faces, makeup transfer on cross-domain faces, video makeup transfer, and makeup-guided text-to-image generation. These results provide compelling evidence of the effectiveness of our approach and highlight its potential for a wide range of applications.
J Broader Impact
The broader impact of diffusion-based makeup transfer methods is significant, as they have the potential to revolutionize the beauty industry by enabling more efficient and personalized makeup application. However, it is important to consider the ethical implications of such technology, including issues related to privacy, consent, and the potential perpetuation of beauty standards. We do not condone the unauthorized use of our method to edit the makeup of others’ photos. As with any emerging technology, we urge caution in the use of diffusion-based makeup transfer methods and further consideration of their ethical and legal implications.
K Limitations
Due to potential inconsistencies in facial structure between paired data in the training dataset, which may arise from conflicts between text editability and maintaining the source image structure in current text-based editing methods, this issue may potentially impact the performance of our model. To address this, we plan to refine the data selection process and manually customize the data. In the future, we also plan to extend our method to 3D tasks.
Figure 11: More results of makeup transfer for realistic faces and cross-domain faces.
Figure 12: More results of video makeup transfer and makeup guided text-to-image generation.

### v2

@@
-https://xiaojiu-z.github.io/Stable-Makeup.github.io/
-Abstract
+Stable-Makeup: When Real-World Makeup Transfer Meets Diffusion Model
+Abstract.
 Current makeup transfer methods are limited to simple makeup styles, making them difficult to apply in real-world scenarios.
@@
 With the aid of our newly added makeup cross-attention layers in U-Net, we can accurately transfer the detailed makeup to the corresponding position in the source image.
-After content-structure decoupling training, Stable-Makeup can maintain content and the facial structure of the source image.
-Moreover, our method has demonstrated strong robustness and generalizability, making it applicable to various tasks such as cross-domain makeup transfer, makeup-guided text-to-image generation and so on.
-Extensive experiments have demonstrated that our approach delivers state-of-the-art (SOTA) results among existing makeup transfer methods and exhibits a highly promising with broad potential applications in various related fields.
-Keywords:
-Makeup Transfer, D-P Makeup Encoder, Diffusion
-Figure 1: Our proposed framework, Stable-Makeup, is a novel diffusion-based method for makeup transfer that can robustly transfer a diverse range of real-world makeup styles, from light to extremely heavy makeup.
-*** Work done during internship at Xiaohongshu Inc.
-$\dagger$ Corresponding author
-1 Introduction
+After content-structure decoupling training, Stable-Makeup can maintain the content and the facial structure of the source image.
+Moreover, our method has demonstrated strong robustness and generalizability, making it applicable to various tasks such as cross-domain makeup transfer, makeup-guided text-to-image generation, and so on.
+Extensive experiments have demonstrated that our approach delivers state-of-the-art results among existing makeup transfer methods and exhibits a highly promising with broad potential applications in various related fields.
+Code released: https://github.com/Xiaojiu-z/Stable-Makeup
+Diffusion Model, Makeup Transfer, Image Generation
+isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
+Figure 1.
+Our proposed framework, Stable-Makeup, is a novel diffusion-based method for makeup transfer that can robustly transfer a diverse range of real-world makeup styles, from light to extremely heavy makeup.
+[…]
+1. Introduction
 As a significant computer vision task, makeup transfer has a wide range of applications, such as in the beauty industry or in virtual try-on systems, and enables the enhancement, modification, and transformation of facial features, achieving the desired effects of beautification, embellishment, and deformation.
 However, despite its conceptual straightforwardness, makeup transfer poses significant challenges when aiming for a seamless and authentic transformation across a wide range of makeup intensities and styles.
-From the perspective of the existing technical route, current makeup transfer techniques [21, 17, 7, 26, 47, 12, 22, 43, 50, 40, 19, 51] rely heavily on Generative Adversarial Network (GAN) based approaches which have demonstrated potential in digital cosmetic applications.
+From the perspective of the existing technical route, current makeup transfer techniques (Li et al., 2018; Jiang et al., 2020; Deng et al., 2021; Nguyen et al., 2021; Xiang et al., 2022; Gu et al., 2019; Liu et al., 2021; Wan et al., 2022; Yan et al., 2023; Sun et al., 2023; Kips et al., 2020; Yang et al., 2022) rely heavily on Generative Adversarial Network (GAN) based approaches which have demonstrated potential in digital cosmetic applications.
 However, these approaches fall short when confronted with the diversity of real-world makeup styles, especially when translating high-detailed and creative cosmetics, such as those found in cosplay or movie character imitations, onto real faces.
@@
 From a data perspective, existing makeup datasets lack diversity and cannot accommodate real-world makeup transfer.
-To address this data shortage, we propose an automatic data construction pipeline that employs large language model and generative model to edit real human face images and create paired before-and-after makeup images.
-Ultimately, our dataset comprises 20k image pairs, covering a wide range of makeup styles from light to heavy makeup.
-This dataset has the potential to facilitate research in the makeup transfer field and enable the development of more robust and accurate models.
-Our Stable-Makeup, which is built on a pre-trained diffusion model, and comprises three key components: the Detail-Preserving Makeup Encoder, Makeup Cross-attention Layers and the Content and Structural Control Modules.
-To preserve makeup details, we employ a multi-layer strategy in our D-P makeup encoder to encode the makeup reference image into multi-scale and spatial-aware detail makeup embeddings.
-The content control module is designed to maintain pixel-level content consistency with source image.
+To address this data shortage, we propose an automatic data construction pipeline that employs a large language model and generative model to edit real human face images and create paired before-and-after makeup images.
+Ultimately, our training data comprises 20k image pairs, covering a wide range of makeup styles from light to heavy makeup.
+This automatic data construction pipeline has the potential to facilitate research in the makeup transfer field and enable the development of more robust and accurate models.
+Our Stable-Makeup, which is built on a pre-trained diffusion model, comprises three key components: the Detail-Preserving Makeup Encoder, Makeup Cross-attention Layers, and the Content and Structural Control Modules.
+To preserve makeup details, we employ a multi-layer strategy in our D-P makeup encoder to encode the makeup reference image into multi-scale detail makeup embeddings.
+The content control module is designed to maintain pixel-level content consistency with the source image.
 The structural control module is utilized to introduce facial structure, improving the consistency between the generated image and the facial structure of the source image.
@@
 In summary, our contributions are:
-  • In this paper, we present Stable-Makeup, the first diffusion-based makeup transfer method to the best of our knowledge.
+  • Stable-Makeup is the first diffusion-based makeup transfer method to the best of our knowledge.
 Our experimental results demonstrate state-of-the-art performance over existing makeup transfer methods.
@@
   • We designed an automatic pipeline to create a high-quality, diverse paired dataset of makeup before and after images.
-After content-structure decoupling training on this dataset, we can further maintain the content and structure of the source image.
-2 Related Works
-2.1 Facial Makeup Transfer
-Traditional makeup transfer methods employ image processing techniques such as facial landmark extraction and detection[49, 41].
-Meanwhile, contemporary deep learning-based makeup transfer methods demonstrate high robustness and generalizability to diverse makeup styles, Generative Adversarial Networks (GANs)[58, 5, 21, 17, 7, 26, 47, 12, 22, 43, 50, 40, 19, 15, 51] as an example, have been extensively applied in facial makeup synthesis tasks.
-In the domain of image transfer, Beauty-GAN[21] driven by pixel-level Histogram Matching and utilizes an array of loss functions to train the primary network.
-PSGAN[17] addresses the challenge of transferring makeup between images with significant facial expression variations and focuses on specific facial regions for transfer.
-To expand the scope of makeup transfer tasks beyond mere color transfer in specific facial areas, CPM[26] introduces pattern addition into the makeup transfer process.
-Furthermore, SCGAN[7] leverages a Part-specific Style Encoder to decouple makeup styles on a component-wise basis.
-Additionally, RamGAN[47] focuses on component consistency by integrating a region-attentive morphing module.
+After content-structure decoupling training on our training data, we can further maintain the content and structure of the source image.
+2. Related Works
+2.1. Facial Makeup Transfer
+Traditional makeup transfer methods employ image processing techniques such as facial landmark extraction and detection(Xu et al., 2013; Tong et al., 2007).
+Meanwhile, contemporary deep learning-based makeup transfer methods demonstrate high robustness and generalizability to diverse makeup styles, Generative Adversarial Networks (GANs)(Zhu et al., 2017; Chang et al., 2018; Li et al., 2018; Jiang et al., 2020; Deng et al., 2021; Nguyen et al., 2021; Xiang et al., 2022; Gu et al., 2019; Liu et al., 2021; Wan et al., 2022; Yan et al., 2023; Sun et al., 2023; Kips et al., 2020; Hu et al., 2022a; Yang et al., 2022) as an example, have been extensively applied in facial makeup synthesis tasks.
+In the domain of image transfer, Beauty-GAN(Li et al., 2018) is driven by pixel-level Histogram Matching and utilizes an array of loss functions to train the primary network.
+However, its effectiveness is limited to frontal facial images, posing a challenge when transferring makeup between images with notable facial pose variations.
+PSGAN(Jiang et al., 2020) addresses the challenge of transferring makeup between images with significant facial expression variations and focuses on specific facial regions for transfer.
+To expand the scope of makeup transfer tasks beyond mere color transfer in specific facial areas, CPM(Nguyen et al., 2021) introduces pattern addition into the makeup transfer process.
+Furthermore, SCGAN(Deng et al., 2021) leverages a Part-specific Style Encoder to decouple makeup styles on a component-wise basis.
+Additionally, RamGAN(Xiang et al., 2022) focuses on component consistency by integrating a region-attentive morphing module.
+Although these GAN-based methods have achieved significant success, they exhibit limitations in transferring complex makeup styles in real-world scenarios.
+Moreover, the majority of these approaches necessitate the alignment of facial regions, further constraining their practical applicability.
 The inadequacies of GAN-based models in addressing makeup transfer problems under real-world makeup settings have necessitated the development of new techniques.
 The primary objective of this paper is to overcome these limitations by utilizing diffusion-based methods to achieve robust and high-quality output for extreme style makeup.
-2.2 Diffusion Models
-Recently, diffusion models have ushered in significant advancements in multi-modal image generation tasks, encompassing text-to-image generation[30, 27, 34, 31, 35], image editing[6, 20, 9, 3, 48, 24, 2, 56, 42], controllable generation[53, 25, 57, 23] and subject-driven generation[55, 14, 1, 16, 10, 33, 32].
-Methods based on diffusion paradigm, along with their subsequent adaptations, have demonstrated state-of-the-art results in various image generation tasks.
-The diffusion model was initially introduced based on the physical inverse diffuse process [37].
-Subsequently, DDPM[13] demonstrated that this approach is viable for image generation by denoising a Gaussian-noised image.
-DDIM[38] introduced a faster non-Markovian sampling process and an implicit modeling of the denoising step, both of which significantly enhance the model’s efficiency.
-Latent Diffusion Models[31] utilize an Autoencoder to encode images into latent embeddings, effectively increasing computational efficiency while preserving the essential features of the input image.
-In recent years, there have been significant advancements in diffusion-based text-to-image generation.
-For instance, Imagen[34] and Stable Diffusion[31] have trained diffusion models on larger datasets, making them the mainstream for generating large images.
-DeepFloyd IF[35] has utilized a triple-cascade diffusion model, which has significantly enhanced the text-to-image generation capability, even generating correct fonts.
-Stable Diffusion XL[27], a two-stage cascade diffusion model, which has greatly improved the generation of high-frequency details, small object features, and overall image color.
-As for controllable generation, ControlNet[54] employs a bypass structure featuring a trainable copy network as the input mechanism for control conditions, effectively integrating a variety of control conditions such as Canny Edge, Depth, and Human Pose, etc.
-As for subject generation, DreamBooth[32] Dreambooth trains a separate Unet for each subject, resulting in excellent consistency of character.
-However, the high computational cost limits its practical application.
-Meanwhile, diffusion based image editing methods have also demonstrated remarkable progress.
-For example, Ledits[42] introduces a training-free image manipulation method, utilizing an inversion approach for processing.
+2.2. Diffusion Models
+Recently, diffusion models have ushered in significant advancements in multi-modal image generation tasks, encompassing text-to-image generation(Ramesh et al., 2022; Podell et al., 2023; Saharia et al., 2022a; Rombach et al., 2022; Saharia et al., 2022b), image editing(Chefer et al., 2023; Li et al., 2023; Epstein et al., 2023; Cao et al., 2023; Xie et al., 2023; Mou et al., 2023a; Bodur et al., 2023; Zhang et al., 2023a; Tsaban and Passos, 2023), controllable generation(Zhang and Agrawala, 2023; Mou et al., 2023b; Zhao et al., 2023; Ma et al., 2023) and subject-driven generation(Zhang et al., 2024; Hu et al., 2022b; Arar et al., 2023; Jia et al., 2023; Gal et al., 2022; Ruiz et al., 2023, 2022).
+Methods based on the diffusion paradigm, along with their subsequent adaptations, have demonstrated state-of-the-art results in various image generation tasks.
+As for controllable generation, ControlNet(Zhang et al., 2023b) employs a bypass structure featuring a trainable copy network as the input mechanism for control conditions, effectively integrating a variety of control conditions such as Canny Edge, Depth, and Human Pose, etc.
+As for subject generation, DreamBooth(Ruiz et al., 2022) Dreambooth trains a separate Unet for each subject, resulting in excellent consistency of character.
+Meanwhile, diffusion-based image editing methods have also demonstrated remarkable progress.
+For example, Ledits(Tsaban and Passos, 2023) introduces a training-free image manipulation method, utilizing an inversion approach for processing.
 Due to the remarkable advancements of diffusion methods in generating high-quality images, benefiting from supervised training on large-scale datasets, they have become the mainstream method in image generation.
@@
 Our approach overcomes the limitations of GAN-based models and achieves robust and high-quality output, particularly for extreme style makeup.
-3 Methodology
-In this part, we introduce the detailed design of our Stable-Makeup in Section 3.1.
-Then, the specifics of our training and inference process are detailed in Section 3.2.
-Preliminaries about diffusion models [31], Controlnet [54] and CLIP [29] can be found in supplementary.
-3.1 Stable-Makeup
-3.1.1 Overview.
+3. Methodology
+In this part, we introduce the detailed design of our Stable-Makeup.
+Then, the specifics of our training and inference process are in a detailed description.
+Preliminaries about diffusion models (Rombach et al., 2022), Controlnet (Zhang et al., 2023b) and CLIP (Radford et al., 2021) can be found in supplementary.
+3.1. Stable-Makeup
+3.1.1. Overview.
 The overall methodology is illustrated in Fig.
@@
 Formally, the target of Stable-Makeup is to transfer the reference makeup from a given image $\mathit{I_{m}}$ onto a user-provided source image $\mathit{I_{s}}$, resulting in a target image $\mathit{I_{t}}$ that exhibits the desired makeup style.
-Therefore, Stable-Makeup necessitates the simultaneous transfer of fine-grained makeup information and preservation of structural information.
-To achieve this, Stable-Makeup first utilizes a Detail-Preserving Makeup Encoder to extract the multi-scale and spatial-aware features of the reference makeup as detailed makeup embeddings $\mathit{E_{m}}$.
-Next, content encoder and structural encoder are employed to encode the source image as content features $\mathit{F_{c}}$ and the facial structure control image $\mathit{I_{c}}$ (obtained from an off-the-shelf face parsing method [28].) as structure features $\mathit{F_{s}}$, respectively.
+Therefore, Stable-Makeup necessitates the simultaneous transfer of fine-grained makeup information and the preservation of structural information.
+To achieve this, Stable-Makeup first utilizes a Detail-Preserving Makeup Encoder to extract the multi-scale features of the reference makeup as detailed makeup embeddings $\mathit{E_{m}}$.
+Next, content encoder and structural encoder are employed to encode the source image as content features $\mathit{F_{c}}$ and the facial structure control image $\mathit{I_{c}}$ (obtained from an off-the-shelf face parsing method (Prados-Torreblanca et al., 2022).) as structure features $\mathit{F_{s}}$, respectively.
 These two features are then incorporated together and fed into the U-Net.
-The detailed makeup embeddings $\mathit{E_{m}}$ along with the content features $\mathit{F_{c}}$ and structure features $\mathit{F_{s}}$ are integrated into the diffusion U-Net as conditions to guide the image generation process.
+The multi-scale detailed makeup embeddings $\mathit{E_{m}}$ along with the content features $\mathit{F_{c}}$ and structure features $\mathit{F_{s}}$ are integrated into the diffusion U-Net as conditions to guide the image generation process.
 During target image $\mathit{I_{t}}$ generation, Stable-Makeup uses newly added cross-attention layers to implicitly align the multi-scale detailed makeup embeddings $\mathit{E_{m}}$ with the intermediate feature maps of facial region in source image at different layers in U-Net.
@@
 With our Content and Structure Decoupling Training, Stable-Makeup can further maintain the content and facial structure of the source image.
-Figure 2: Overall schematics of our method.
-Given a source image $\mathit{I_{s}}$, a reference makeup image $\mathit{I_{m}}$ and an obtained facial structure control image $\mathit{I_{c}}$, Stable-Makeup utilizes D-P makeup encoder to encode $\mathit{I_{m}}$.
+Figure 2.
+Overall schematics of our method.
+Given a source image $\mathit{I_{s}}$, a reference makeup image $\mathit{I_{m}}$ and an obtained facial structure control image $\mathit{I_{c}}$, Stable-Makeup utilizes the D-P makeup encoder to encode $\mathit{I_{m}}$.
 Content and structural encoders are used to encode $\mathit{I_{s}}$ and $\mathit{I_{c}}$ respectively.
-With the aid of the makeup cross-attention layers, Stable-Makeup aligns the facial regions of $\mathit{I_{s}}$ and $\mathit{I_{m}}$, enabling successful transfers the intricate makeup details.
+With the aid of the makeup cross-attention layers, Stable-Makeup aligns the facial regions of $\mathit{I_{s}}$ and $\mathit{I_{m}}$, enabling successful transfers of the intricate makeup details.
 After content-structure decoupling training, Stable-Makeup further maintains content and structure of $\mathit{I_{s}}$.
-3.1.2 Detail-Preserving Makeup Encoder.
+3.1.2. Detail-Preserving Makeup Encoder.
 Makeup images often involve intricate details such as eyelashes, eyebrows, and fine lines.
@@
 […]
-Notably, some previous works[52, 46, 11] utilize linear layer mapping of CLIP features.
+Notably, some previous works(Ye et al., 2023; Wei et al., 2023; Gal et al., 2023) utilize linear layer mapping of CLIP features.
 This design cannot preserve the spatial information between input features.
 Different from them, we utilize a self-attention layer to map the multi-layer CLIP embeddings (256-dimensional patch features and 1-dimensional CLS features) of the reference makeup image and concatenate on feature dimension.
-This approach effectively preserves the spatial, multi-scale, and global information of the makeup image, thereby facilitating improved alignment of intermediate related features within the U-Net architecture.
+This approach effectively preserves the spatial, multi-scale, and global information of the makeup image, thereby facilitating improved alignment of intermediate-related features within the U-Net architecture.
 Consequently, it yields more precise preservation of fine-grained details and enhanced transfer outcomes.
@@
 Furthermore, we provide additional comparative analysis in supplementary to demonstrate the superiority of our method.
-3.1.3 Makeup Cross-attention Layers.
-The goal of the makeup cross-attention layers is to integrate multi-scale detailed makeup embeddings into the U-Net and align them appropriately with the intermediate feature maps of facial region in the source image, enabling the model to accurately transfer the makeup onto the facial region in the source image.
+3.1.3. Makeup Cross-attention Layers.
+The goal of the makeup cross-attention layers is to integrate multi-scale detailed makeup embeddings into the U-Net and align them appropriately with the intermediate feature maps ofthe facial region in the source image, enabling the model to accurately transfer the makeup onto the facial region in the source image.
 Specifically, as shown in the figure 2, in each transformer block of the U-Net, we retain the original text cross-attention and add Makeup Cross-attention Layers.
-The multi-scale detailed makeup embeddings $\mathit{E_{m}}$ are input to the makeup cross-attention layers, serves as the K(key) and V(value) features.
+The multi-scale detailed makeup embeddings $\mathit{E_{m}}$ are input to the makeup cross-attention layers and serve as the K(key) and V(value) features.
 The makeup cross-attention layers and the text cross-attention layers in the original U-Net share the Q feature.
-With the help of the cross-attention mechanism, the intermediate facial features of source image and makeup image are well-aligned in the stable diffusion U-Net.
+With the help of the cross-attention mechanism, the intermediate facial features of the source image and makeup image are well-aligned in the stable diffusion U-Net.
 Finally, we simply add the output of makeup cross-attention to the output of text cross-attention.
@@
 In this task, the reference makeup image is employed as a guiding condition in conjunction with the original text prompt to influence the pre-trained U-Net architecture, and the makeup consistency is maintained during the text-to-image generation process.
-3.1.4 Content and Structural Control Modules.
-To maintain the structural consistency of the source image during makeup transfer, we utilize Content and Structural Control Modules, which consists of two adapted ControlNets serving as encoders.
-The target of the content encoder is to maintain content consistency and consistency in non-facial areas of the original image, and it uses the ControlNet to encode the source image.
+3.1.4. Content and Structural Control Modules.
+To maintain the structural consistency of the source image during makeup transfer, we utilize Content and Structural Control Modules, which consist of two adapted ControlNets serving as encoders.
+The target of the content encoder is to maintain content consistency and consistency in non-facial areas of the original image, and it uses ControlNet to encode the source image.
 To control facial structures during the makeup transfer process, we added an extra structural encoder.
@@
 Finally, we directly add the outputs of the content encoder and structural encoder to the Stable Diffusion model.
-3.2 Model Training and Inference
-3.2.1 Makeup-real Dataset.
+3.2. Model Training and Inference
+3.2.1. Training data collection.
 As depicted in Fig.
-3, Ours pipeline to create a large-scale makeup pairing dataset (Makeup-real dataset) by leveraging the capabilities of LLM and pre-trained diffusion involves three main steps: Firstly, we utilize GPT4 to generate 1000 prompts using the template "Make it {} makeup" to cover a wide range of makeup styles.
+3, our pipeline to create a large-scale makeup pairing dataset by leveraging the capabilities of LLM and pre-trained diffusion involves three main steps: Firstly, we utilize GPT4 to generate 1000 prompts using the template “Make it {} makeup” to cover a wide range of makeup styles.
 After filtering out duplicate words, we obtain 300 different prompts.
-Secondly, we sample 20,000 faces from the FFHQ[18] dataset as non-makeup images and apply makeup to these images using existing text-based real image editing algorithm [42].
+Secondly, we sample 20,000 faces from the FFHQ(Karras et al., 2019) dataset as non-makeup images and apply makeup to these images using existing text-based real image editing algorithm (Tsaban and Passos, 2023).
 For each face image, we randomly select a makeup style from the 300 prompts obtained in the first step, resulting in 20,000 face images with consistent identities and makeup styles.
-Finally, to prevent the unwanted effects of editing non-face regions during the edit process, we use a face segmentation model to separate the face region from the background and other non-face regions.
+Finally, to prevent the unwanted effects of editing non-face regions during the editing process, we use a face segmentation model to separate the face region from the background and other non-face regions.
 We composite the non-face regions of the original face images onto the edited face images to obtain the final makeup pairing dataset.
-[Uncaptioned image] Figure 3: The procedure for training data generation.
-Finally, our dataset includes 20,000 image pairs covering a wide range of makeup styles from light to heavy makeup, represents a significant improvement over previous datasets in terms of diversity and quantity.
+Figure 3. The procedure for training data generation.
+Finally, our training data includes 20,000 image pairs covering a wide range of makeup styles from light to heavy makeup, which represents a significant improvement over previous datasets in terms of diversity and quantity.
 In our supplementary materials, we also present additional visual results for the alternative methods in Step 2.
-3.2.2 Content-Structure Decoupling Training.
+Figure 4. Qualitative comparison of different methods. Our results outperform other methods in terms of makeup detail transfer, ranging from light makeup to personalized heavy makeup.
+3.2.2. Content-Structure Decoupling Training.
 The current diffusion-based real image editing faces a conflict between maintaining the original layout and editability.
@@
 By decoupling content and structure during training, we can control and adjust the content and structure separately, improving the accuracy and consistency of the generated images.
-During inference, $\mathit{I_{s}}$ and $\mathit{I_{c}}$ obtained from the source image are input to content encoder and structural encoder respectively.
-3.2.3 Augmentation.
+During inference, $\mathit{I_{s}}$ and $\mathit{I_{c}}$ obtained from the source image are input to the content encoder and structural encoder respectively.
+3.2.3. Augmentation.
 In our training pipeline, we have used a variety of augmentations, which are crucial for adapting to real-world scenarios and achieving successful makeup transfer.
 These augmentations can be divided into two categories: makeup augmentations and source augmentations.
-Makeup augmentations involves structural augmentations such as face warping and affine transformations applied to the input of the makeup encoder.
+Makeup augmentations involve structural augmentations such as face warping and affine transformations applied to the input of the makeup encoder.
 The purpose of these augmentations is to eliminate Content information from the reference makeup image and to disrupt the pixel alignment between the reference makeup image and the source image.
@@
 These augmentations include synchronized affine transformations applied to both the source image and the target image.
-3.2.4 Loss function.
+3.2.4. Loss function.
 Our loss function is similar to the original Stable Diffusion training objective function and can be mathematically represented as:
-$L(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\mathbf{c_{i}},\mathbf{c_{e}},\mathbf{c_{m}}\right)\right\|_{2}^{2}\right],$ (1)
+(1) $L(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\mathbf{c_{i}},\mathbf{c_{e}},\mathbf{c_{m}}\right)\right\|_{2}^{2}\right],$
 […]
-4 Experiments
-4.1 Experimental Setup
-We employed Stable Diffusion V1-5 as the pre-trained diffusion model and trained it on the Makeup-real Dataset, which consists of 20k makeup-non-makeup pairs and each pair with the same identity.
-We focus on real-world conditions and utilize extremely aggressive augmentation, as described in Section 3.2.
-While training, the parameters of the D-P makeup encoder, makeup cross-attention layers, content encoder and structural encoder are updated, while the pre-trained text-to-image model remains frozen.
+Figure 5. Qualitative ablation results. Figure (a) explores the classifier-free guidance scale. Figure (b) presents the ablation study of different training and inference settings.
+4. Experiments
+Method CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$
+(M-wild dataset) (CPM-real dataset)
+BeautyGAN 0.742 0.926 0.884 48.408 0.718 0.915 0.866 49.168
+CPM 0.754 0.920 0.873 94.092 0.757 0.920 0.846 95.172
+PSGAN 0.748 0.928 0.862 45.742 0.724 0.916 0.829 46.091
+EleGANt 0.759 0.928 0.878 44.463 0.731 0.917 0.838 45.797
+SCGAN 0.751 0.928 0.882 44.296 0.726 0.916 0.858 46.284
+SPMT 0.761 0.931 0.915 29.126 0.735 0.920 0.900 28.299
+SSAT 0.758 0.930 0.914 32.578 0.719 0.915 0.896 33.604
+Ours 0.768 0.933 0.910 29.087 0.760 0.923 0.868 27.976
+Table 1. Quantitative comparison of different methods. Metrics that are red and blue represent methods that rank 1st and 2nd, respectively.
+4.1. Experimental Setup
+We employed Stable Diffusion V1-5 as the pre-trained diffusion model and trained it on the Makeup-real Dataset, which consists of 20k makeup-non-makeup pairs, and each pair with the same identity.
+We focus on real-world conditions and utilize extremely aggressive augmentation.
+While training, the parameters of the D-P makeup encoder, makeup cross-attention layers, content encoder, and structural encoder are updated, while the pre-trained text-to-image model remains frozen.
 The model underwent 100,000 iterations of training on 8 H800 GPUs, with a batch size of 16 per GPU and a learning rate of 5e-5.
 Inference was performed using DDIM as the sampler, with a step size of 30 and a guidance scale set to 1.5.
-4.2 Evaluation Benchmarks
-  • CPM-real[26] dataset: It is very diverse in terms of makeup styles, containing both color and pattern makeups.
-The degree of makeup can vary from light to heavy, from color-oriented to pattern-driven.
-  • Makeup-wild[17] dataset: It contains facial images with various poses and expressions as well as complex backgrounds to test methods in the real-world environment.
-4.3 Evaluation Metrics
-To provide a thorough and objective assessment of the performance of each algorithm in different aspects of makeup transfer, we employed the following metrics: CLIP-I[29] and DINO-I[4] were utilized to assess makeup transfer performance.
+4.2. Evaluation Metrics
+To provide a thorough and objective assessment of the performance of each algorithm in different aspects of makeup transfer, we employed the following metrics: CLIP-I(Radford et al., 2021) and DINO-I(Caron et al., 2021) were utilized to assess makeup transfer performance.
 After makeup transfer, the content and structure information of the input image should not be changed.
-So we use SSIM [45] and L2-M to evaluate the degree of content and structural maintenance.
+So we use SSIM (Wang et al., 2004) and L2-M to evaluate the degree of content and structural maintenance.
 Notably, CLIP-I and DINO-I are metrics used to calculate the cosine similarity between the image embeddings of the target image and the makeup image.
@@
 L2-M metric utilizes the facial segmentation method to obtain the background images of the target and the source image, calculate the square of the difference, and then average it based on the number of pixels.
-4.4 Comparison Methods
-We performed comprehensive comparisons with the most representative makeup transfer algorithms currently available, including BeautyGAN[21], CPM[26], PSGAN[17], EleGANt[51], SCGAN[7], SPMT[59] and SSAT[40], and conduct comparisons on M-Wild dataset and CPM-real dataset respectively.
-4.5 Experiment Results
-4.5.1 Qualitative Comparison.
-Figure 4: Qualitative comparison of different methods. Our results outperform other methods in terms of makeup detail transfer, ranging from light makeup to personalized heavy makeup.
+4.3. Comparison Methods
+We performed comprehensive comparisons with the most representative makeup transfer algorithms currently available, including BeautyGAN(Li et al., 2018), CPM(Nguyen et al., 2021), PSGAN(Jiang et al., 2020), EleGANt(Yang et al., 2022), SCGAN(Deng et al., 2021), SPMT(Zhu et al., 2022) and SSAT(Sun et al., 2023), and conduct comparisons on M-Wild dataset(Jiang et al., 2020) and CPM-real dataset(Nguyen et al., 2021) respectively.
+4.4. Experiment Results
+4.4.1. Qualitative Comparison.
+Figure 6. The visual results of Stable-Makeup demonstrate its robustness and generalizability for various applications.
 Fig.
@@
 In terms of makeup transfer capability, our method outperformed other methods by a significant margin.
-Other methods were only effective in transferring simple lip and eye shadow colors, and were only effective for light makeup.
-In contrast, our method not only performed well for light makeup, but also effectively aligned and transferred heavy makeup to the corresponding positions in the source image, while preserving makeup details and corresponding color distributions on the facial regions of the makeup image.
-4.5.2 Quantitative Comparison.
-Table 1: Quantitative comparison of different methods. Metrics that are bold and underlined represent methods that rank 1st and 2nd, respectively.
-Method CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$ CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$
-(M-wild dataset) (CPM-real dataset)
-BeautyGAN 0.818 0.927 0.878 48.775 0.775 0.913 0.858 53.323
-CPM 0.815 0.923 0.862 108.374 0.793 0.921 0.837 114.288
-PSGAN 0.739 0.929 0.850 48.816 0.693 0.915 0.839 50.137
-EleGANt 0.827 0.928 0.891 46.939 0.787 0.917 0.872 48.995
-SCGAN 0.816 0.927 0.878 47.595 0.776 0.915 0.851 50.062
-SPMT 0.834 0.930 0.911 31.671 0.778 0.918 0.905 32.323
-SSAT 0.823 0.931 0.901 34.512 0.773 0.910 0.881 35.205
-Ours 0.860 0.935 0.906 31.483 0.825 0.928 0.883 32.255
-To test the ability of methods in the real-world environment, We use the non-makeup 334 images of the Makeup-Wild(M-Wild) dataset as source images, and randomly selected 334 images from the makeup images in the M-Wild dataset and the CPM-real dataset, respectively, as our reference makeup images.
+Other methods were only effective in transferring simple lip and eye shadow colors and were only effective for light makeup.
+In contrast, our method not only performed well for light makeup but also effectively aligned and transferred heavy makeup to the corresponding positions in the source image, while preserving makeup details and corresponding color distributions on the facial regions of the makeup image.
+4.4.2. Quantitative Comparison.
+To test the ability of methods in the real-world environment, We utilized all makeup images from M-wild Dataset and CPM-Real dataset as reference images, while randomly selecting non-makeup images from the MT-dataset and MWild-Dataset as source images.
+We utilized a total of 7017 makeup reference images, paired with an equal number of source images.
 As shown in Table 1, It is evident that Stable-Makeup outperforms all other makeup transfer methods in terms of the DINO-I and CLIP-I metrics, indicating that our method has superior makeup transfer capabilities compared to other methods.
@@
 It should be noted that SSIM metrics may not accurately reflect the makeup transfer performance, as they may produce good scores when the transferred result is identical to the source image, indicating that the transfer algorithm did not work.
-In line with previous studies[26, 47, 59], we also listed it as a reference in our evaluation.
-4.5.3 User Study.
-Table 2: User study. Metrics that are bold represent method that rank 1st. Our method achieves the highest average score.
-Method BeautyGAN CPM PSGAN EleGANt SCGAN SPMT SSAT Ours
-Average Score 3.01 3.13 2.95 3.62 2.90 3.35 2.85 4.58
-We conducted a user study to compare our method with other methods perceptually.
-We randomly select 10 source images and 10 reference images from both the M-Wild dataset and CPM-real dataset.
-For each evaluation, each user will see one source image with a reference makeup image, and one image generated by each method.
-60 evaluators were asked to provide a comprehensive score for each generated image based on its makeup consistency, content and structure consistency, using a scale of 1 (worst) to 5 (best).
-The results of average scores in Table 2 indicate that our method outperforms the comparison methods.
-Our method achieves the highest average score, which clearly indicates that our method outperforms the comparison methods.
-4.5.4 Ablation Study.
-Figure 5: Qualitative ablation results. Figure (a) explores the classifier-free guidance scale. Figure (b) presents the ablation study of different training and inference settings.
-Ablation Setups CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$
-(CPM-real dataset)
-W.O. structure input 0.776 0.923 0.772 103.029
-W.O. source aug 0.816 0.921 0.874 153.463
-W.O decoupling training 0.806 0.927 0.791 32.856
-W.O multi-layer 0.803 0.911 0.880 34.861
-Ours full 0.825 0.928 0.883 32.255
-Table 3: Quantitative ablation results on CPM-real dataset.
+In line with previous studies(Nguyen et al., 2021; Xiang et al., 2022; Zhu et al., 2022), we also listed it as a reference in our evaluation.
+4.4.3. Ablation Study.
 Firstly, we explored the effect of different parameters for the classifier-free guidance scale (GS) under makeup conditions in figure 5 (a).
@@
 Removing source augmentation during training (third column), the alignment and transferability of the model were weakened, and the background color was changed.
-When removing structural encoder during training (fourth column, W.O.
+When removing the structural encoder during training (fourth column, W.O.
 decoupling training), the facial structure also could not be well preserved.
@@
 When our method was fully implemented, it achieved a complete transfer of makeup details while maintaining good consistency of the content and structure of the source image.
-Quantitative ablation results are shown in Table 3.
-We performed quantitative ablation experiments using the same experimental setup on CPM-real dataset as Sec.
-4.5.
-We observed that our method achieved optimal performance when all modules were retained, while removing the structural encoder input during inference (first row) or training (third row) led to a significant decrease in content and structure consistency.
-Removing the multi-layer strategy during training resulted in a significant decrease in makeup transfer ability, and the absence of source enhancement led to a significant decrease in background consistency.
-We exclude the setting without makeup augmentations, as shown in Figure 5 (b) second column, where the model can only perform face swapping and lacks the ability to perform makeup transfer.
-This could significantly affect the calculation of our makeup transfer metrics.
-Figure 6: The visual results of Stable-Makeup demonstrate its robustness and generalizability for various applications.
-4.6 More Applications
-As illustrated in the Fig.
-6, our Stable-Makeup not only achieves unprecedented makeup transfer effects on real faces, but also enables various applications that were previously unattainable with traditional makeup transfer methods.
+Furthermore, we provide additional quantitative ablation results in supplementary.
+4.5. More Applications
+As illustrated in Fig.
+6, our Stable-Makeup not only achieves unprecedented makeup transfer effects on real faces but also enables various applications that were previously unattainable with traditional makeup transfer methods.
 For instance, our method can perform makeup transfer on cross-domain faces and guide text-to-image generation with reference makeup conditions to create creative artistic images.
@@
 providing a broader range of creative makeup options for real human faces.
-5 Conclusion
-In this paper, we propose the Stable-Makeup, a diffusion-based framework aiming for real-world makeup transfer.
+5. Conclusion
+In this paper, we propose the Stable-Makeup, a novel diffusion-based framework aiming for real-world makeup transfer.
 This method has made a significant breakthrough in the field of makeup transfer by achieving effects that were previously unattainable.
-The Stable-Makeup consists of detail-preserving makeup encoder, makeup cross-attention layers, content and structural control modules.
-The detail-preserving makeup encoder captures fine details of makeup, while two encoder structures are employed for the content and structural control modules to encode content and structure of source image.
-The makeup cross-attention layers effectively aligns facial region between reference makeup and source image.
-Through the content and structure decoupling training, we are able to further maintain content and structure consistency.
-In addition, we proposed a automatic pipeline to create a variety of makeup pairing data for training.
+The Stable-Makeup consists of three core modules.
+The detail-preserving makeup encoder captures intricate makeup details, integrating with two modules for content and structural control of the source image.
+Makeup cross-attention layers align facial regions between reference and source images, while content and structure decoupling training enhances consistency.
+Additionally, we propose an automated pipeline for generating diverse makeup pairing data for training.
 Our proposed method has surpassed all existing makeup transfer algorithms and has expanded the boundaries of the makeup transfer field.
-This breakthrough has significant implications for the practical application of makeup transfer technology in various fields, such as cosmetics, entertainment, and fashion.
+This breakthrough has significant implications for the practical application of makeup transfer technology in various fields.
 Supplementary Materials
-F Preliminaries
-F.0.1 Diffusion models.
-Diffusion Model (DM) [13, 39] belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to target data distribution $\mathbf{x_{0}}$ by means of an iterative denoising procedure.
-Latent Diffusion Model (LDM) [31] is proposed to model image representations in autoencoder’s latent space.
+Appendix A Preliminaries
+A.0.1. Diffusion models.
+Diffusion Model (DM) (Ho et al., 2020; Song et al., 2020) belongs to the category of generative models that denoise from a Gaussian prior $\mathbf{x_{T}}$ to target data distribution $\mathbf{x_{0}}$ by means of an iterative denoising procedure.
+Latent Diffusion Model (LDM) (Rombach et al., 2022) is proposed to model image representations in the autoencoder’s latent space.
 LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions.
 The LDM loss is:
-$L_{LDM}(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\bm{\tau_{\theta}}(\mathbf{c})\right)\right\|_{2}^{2}\right],$ (2)
+(2) $L_{LDM}(\bm{\theta}):=\mathbb{E}_{\mathbf{x_{0}},t,\bm{\epsilon}}\left[\left\|\bm{\epsilon}-\bm{\epsilon_{\theta}}\left(\mathbf{x_{t}},t,\bm{\tau_{\theta}}(\mathbf{c})\right)\right\|_{2}^{2}\right],$
 […]
 Stable Diffusion (SD) is a widely adopted text-to-image diffusion model based on LDM.
-Compared to LDM, SD is trained on a large LAION [36] dataset and replaces BERT with the pre-trained CLIP [29] text encoder.
-F.0.2 ControlNet.
-ControlNet [54] is one of the most popular control modules in current diffusion models.
+Compared to LDM, SD is trained on a large LAION (Schuhmann et al., 2022) dataset and replaces BERT with the pre-trained CLIP (Radford et al., 2021) text encoder.
+A.0.2. ControlNet.
+ControlNet (Zhang et al., 2023b) is one of the most popular control modules in current diffusion models.
 It receives inputs from various modalities as spatial control conditions, guiding the diffusion model to generate desired images, thereby achieving controllable generation.
@@
 The complete ControlNet can represented as:
-$\boldsymbol{y}_{\mathrm{c}}=\mathcal{F}(\boldsymbol{x};\Theta)+\mathcal{Z}\left(\mathcal{F}\left(\boldsymbol{x}+\mathcal{Z}\left(\boldsymbol{c};\Theta_{\mathrm{z}1}\right);\Theta_{\mathrm{c}}\right);\Theta_{\mathrm{z}2}\right)$ (3)
+(3) $\bm{y}_{\mathrm{c}}=\mathcal{F}(\bm{x};\Theta)+\mathcal{Z}\left(\mathcal{F}\left(\bm{x}+\mathcal{Z}\left(\bm{c};\Theta_{\mathrm{z}1}\right);\Theta_{\mathrm{c}}\right);\Theta_{\mathrm{z}2}\right)$
 Where $\Theta_{\mathrm{z}1}$ and $\Theta_{\mathrm{z}2}$ represent two different zero conv layers’ parameters respectively.
-$\mathcal{F}$ represent diffusion U-Net, $\boldsymbol{x}$ represent image latent, $\Theta$ is the frozen weight of the U-Net and $\Theta_{\mathrm{c}}$ is the trainable copy weight of the U-Net.
-F.0.3 CLIP.
-CLIP [29] consists of two integral components: an image encoder, represented as $F(x)$, and a text encoder, represented as $G(t)$.
+$\mathcal{F}$ represent diffusion U-Net, $\bm{x}$ represent image latent, $\Theta$ is the frozen weight of the U-Net and $\Theta_{\mathrm{c}}$ is the trainable copy weight of the U-Net.
+A.0.3. CLIP.
+CLIP (Radford et al., 2021) consists of two integral components: an image encoder, represented as $F(x)$, and a text encoder, represented as $G(t)$.
 The image encoder, $F(x)$, transforms an image $x$ with dimensions $\mathbb{R}^{3\times H\times W}$ (height $H$ and width $W$) into a $d$-dimensional image feature $f_{x}$ with dimensions $\mathbb{R}^{N\times d}$, where $N$ is the number of divided patches.
@@
 After training on a contrastive loss function, CLIP can be applied directly for zero-shot image recognition without the need for fine-tuning the entire model.
-G Further Exploration of Data Construction
+Appendix B Evaluation Benchmarks
+  • CPM-real dataset: It is very diverse in terms of makeup styles, containing both color and pattern makeup. The degree of makeup can vary from light to heavy, from color-oriented to pattern-driven.
+  • M-wild dataset: It contains facial images with various poses and expressions as well as complex backgrounds to test methods in a real-world environment.
+Ablation Setups CLIP-I $\uparrow$ DINO-I $\uparrow$ SSIM $\uparrow$ L2-M $\downarrow$
+(CPM-real dataset)
+W.O. structure input 0.716 0.913 0.759 97.746
+W.O. source aug 0.752 0.911 0.858 96.385
+W.O decoupling training 0.745 0.917 0.754 35.675
+W.O multi-layer 0.742 0.907 0.861 29.076
+Ours full 0.760 0.923 0.868 27.976
+Table 2. Quantitative ablation results on CPM-real dataset.
+Appendix C Quantitative Ablation
+Quantitative ablation results are shown in Table 2.
+We performed quantitative ablation experiments using the same experimental setup on the CPM-real dataset as the experiments section.
+We observed that our method achieved optimal performance when all modules were retained while removing the structural encoder input during inference (first row) or training (third row) led to a significant decrease in content and structure consistency.
+Removing the multi-layer strategy during training resulted in a significant decrease in makeup transfer ability, and the absence of source enhancement led to a significant decrease in background consistency.
+We exclude the setting without makeup augmentations, as shown in Figure 5 (b) second column in our paper, where the model can only perform face swapping and lacks the ability to perform makeup transfer.
+This could significantly affect the calculation of our makeup transfer metrics.
+Appendix D Further Exploration of Data Construction
 Currently, diffusion-based methods have demonstrated remarkable performance in terms of fidelity and diversity, making them a crucial tool for achieving high-quality and diverse data in step 2 of our data construction process.
-In addition to the inversion-based image editing methods we employed, we could also choose to use existing identity customization methods, such as InstantID[44], which do not require test-time fine-tuning.
+In addition to the inversion-based image editing methods we employed, we could also choose to use existing identity customization methods, such as InstantID(Wang et al., 2024), which do not require test-time fine-tuning.
 However, as shown in figure 7, when using these diffusion-based methods for editing facial images, visible changes in facial structure may occur during makeup generation guided by text.
 For example, the mouth shape of these generated images cannot be exactly the same.
-Given the current limitations of diffusion-based methods in terms of fine-grained control, our proposed content and structure decoupling strategy is particularly important for maintaining the facial expression and structure.
-Figure 7: Visual results of different data construct pipeline.
-H Designing Choice of Detail-Preserving Makeup Encoder
-In this section, we conducted a comparative study between different mapping structure setting under the same training conditions to demonstrate that mapping without preserving spatial information struggling to align and preserving makeup details at a fine-grained level, and to validate the effectiveness of our proposed method.
+Given the current limitations of diffusion-based methods in terms of fine-grained control, our proposed content and structure decoupling strategy is particularly important for maintaining facial expression and structure.
+Figure 7. Visual results of different data construct pipelines.
+Appendix E Designing Choice of Detail-Preserving Makeup Encoder
+In this section, we conducted a comparative study between different mapping structure settings under the same training conditions to demonstrate that mapping without preserving spatial information struggles to align and preserve makeup details at a fine-grained level, and to validate the effectiveness of our proposed method.
 We first introduced our experimental settings in Section 8.
-Subsequently, we conduct visual comparison in Sec.
-H.2 to demonstrate the pivotal role of our proposed method in preserving and aligning the makeup details.
+Subsequently, we conduct a visual comparison in Sec.
+E.2 to demonstrate the pivotal role of our proposed method in preserving and aligning the makeup details.
 Additionally, in Sec.
-H.3 we further demonstrated the effectiveness of our proposed method in preserving and aligning the makeup details by visualizing the attention maps of the makeup cross-attention layers in U-Net.
-H.1 Experimental Setup
-Figure 8: The structural composition of the five distinct mapping schemes. We omitted the drawing of the global vector (cls token from CLIP) for the sake of brevity and clarity.
+E.3 we further demonstrated the effectiveness of our proposed method in preserving and aligning the makeup details by visualizing the attention maps of the makeup cross-attention layers in U-Net.
+E.1. Experimental Setup
+Figure 8. The structural composition of the five distinct mapping schemes. We omitted the drawing of the global vector (cls token from CLIP) for the sake of brevity and clarity.
 As shown in Fig. 8, we conducted training on five sets of mapping schemes.
@@
   • Spatial linear mapping: We utilized linear layers to project the feature dimension to a smaller dimension without affecting the spatial features dimension, where the original dimension was (1, 257$\times$12, 1024), and the projected dimension was (1, 257$\times$12, 768).
-  • Self-Attention mapping: We utilized our proposed self-attention mapping in section 3.1.2. The projected dimension is the same as the original dimension (1, 257$\times$12, 1024).
+  • Self-Attention mapping: We utilized our proposed self-attention mapping in the Methodology section. The projected dimension is the same as the original dimension (1, 257$\times$12, 1024).
   • IP-Adapter: We replaced our D-P makeup encoder with an IP-Adapter that projects the pooled feature of CLIP (1, 1, 768) to a higher dimension features (1, 4, 768) using linear layer and layer norm.
@@
 It is noteworthy that only spatial linear mapping and our self-attention mapping preserve the spatial information of the makeup image features among the five mapping settings, while the others introduce varying degrees of spatial distortion to the makeup image features.
-Figure 9: Visual comparison results under the same training step but with different mapping settings.
-H.2 Comparison
+Figure 9. Results under extreme viewing angles. (please zoom in to better view).
+Figure 10. Visual comparison results under the same training step but with different mapping settings.
+E.2. Comparison
 Under the same training conditions, the makeup transfer results at different training steps are presented in figure 10.
@@
 These experimental results effectively demonstrate the importance of preserving spatial information for preserving fine makeup details.
-H.3 Visualization of Attention maps
-Figure 10: Visualization of attention maps with different mapping settings at the 50,000th training step.
+E.3. Visualization of Attention maps
+Figure 11. Visualization of attention maps with different mapping settings at the 50,000th training step.
 In order to conduct a more comprehensive comparison between spatial linear mapping and self-attention mapping, we visualized the attention maps of the makeup cross-attention layers under both configurations at the 50,000th training step in this experiment.
 Specifically, we extract and aggregate the attention matrices calculated from the patch features of the source facial region (Query) and the multi-scale makeup embeddings (Key-Value), based on the index of the facial position, and visualize them at different layers, as shown in Figure 11 (a).
-From the attention maps, we can see that the self-attention layer can effectively weight the features, resulting in high response for the facial patch features in between the source image and the makeup image, while low response between the other positions in the makeup image.
-However, when we replace the self-attention mapping with spatial linear mapping, the facial region of the source image also exhibit high response to the background features, as shown in the attention maps of layer1-4 in Figure 10 (b).
+From the attention maps, we can see that the self-attention layer can effectively weigh the features, resulting in a high response for the facial patch features in between the source image and the makeup image, while a low response between the other positions in the makeup image.
+However, when we replace the self-attention mapping with spatial linear mapping, the facial region of the source image also exhibits a high response to the background features, as shown in the attention maps of layers 1-4 in Figure 11 (b).
 These findings highlight the advantages of self-attention mapping in achieving more precise and effective feature weighting, which is crucial for achieving high-quality makeup transfer results.
-I More Results
+Appendix F More Results
 In this section, we present a range of additional results that demonstrate the robustness and superiority of our approach.
@@
 12 and Fig.
-12, we showcase the effectiveness of our method in makeup transfer on realistic faces, makeup transfer on cross-domain faces, video makeup transfer, and makeup-guided text-to-image generation.
+13, we showcase the effectiveness of our method in makeup transfer on realistic faces, makeup transfer on cross-domain faces, video makeup transfer, and makeup-guided text-to-image generation.
 These results provide compelling evidence of the effectiveness of our approach and highlight its potential for a wide range of applications.
-J Broader Impact
-The broader impact of diffusion-based makeup transfer methods is significant, as they have the potential to revolutionize the beauty industry by enabling more efficient and personalized makeup application.
+Appendix G User study
+Table 3. User study. Metrics that are bold represent methods that rank 1st. Our method achieves the highest average score.
+Method BeautyGAN CPM PSGAN EleGANt SCGAN SPMT SSAT Ours
+Average Score 3.01 3.13 2.95 3.62 2.90 3.35 2.85 4.58
+We conducted a user study to compare our method with other methods perceptually.
+We randomly select 10 source images and 10 reference images from both the M-Wild dataset and the CPM-real dataset.
+For each evaluation, each user will see one source image with a reference makeup image and one image generated by each method.
+60 evaluators were asked to provide a comprehensive score for each generated image based on its makeup consistency, content and structure consistency, using a scale of 1 (worst) to 5 (best).
+The results of average scores in Table 3 indicate that our method outperforms the comparison methods.
+Our method achieves the highest average score, which clearly indicates that our method outperforms the comparison methods.
+Appendix H Extreme viewing angles
+Our approach is capable of processing source images under extreme viewing angles (in Fig. 9).
+Appendix I Broader Impact
+The broader impact of diffusion-based makeup transfer methods is significant, as they have the potential to revolutionize the beauty industry by enabling more efficient and personalized makeup applications.
 However, it is important to consider the ethical implications of such technology, including issues related to privacy, consent, and the potential perpetuation of beauty standards.
@@
 As with any emerging technology, we urge caution in the use of diffusion-based makeup transfer methods and further consideration of their ethical and legal implications.
-K Limitations
+Appendix J Limitations
 Due to potential inconsistencies in facial structure between paired data in the training dataset, which may arise from conflicts between text editability and maintaining the source image structure in current text-based editing methods, this issue may potentially impact the performance of our model.
@@
 In the future, we also plan to extend our method to 3D tasks.
-Figure 11: More results of makeup transfer for realistic faces and cross-domain faces.
-Figure 12: More results of video makeup transfer and makeup guided text-to-image generation.
+Figure 12. More results of makeup transfer for realistic faces and cross-domain faces.
+Figure 13. More results of video makeup transfer and makeup guided text-to-image generation.